i have to find an equation of a line that goes through the point (2,1) and has a slope of 4.
i have to leave it in slope intercept form.

Answer by scott8148(6628)

start by writing the equation in point-slope form ... (y-1)=4(x-2) ... then convert to slope-intercept

y-1=4x-8 ... y=4x-7